Hi there,
I am trying to find an example for the folowing equation:
sqrt( a^2+b^2)≠a+b
Your help is greatly appreciated.
Joanna

Answer by jim_thompson5910(35256)

Ex: Let a = 3 and b = 4 (pick any two random numbers, just make sure that neither number is 0)


sqrt( a^2+b^2) = a+b

sqrt( 3^2+4^2) = 3+4

sqrt( 9+16) = 3+4

sqrt(25) = 3+4

5 = 3+4

5 = 7

5 ≠ 7


So sqrt( a^2+b^2)≠a+b in general.